OpenStudy (this):

some one help me plllzzz factor by grouping;leading coefficient >1 20m^3+28^2+49

OpenStudy (anonymous):

Is it 28m not 28?

OpenStudy (this):

yhhh right 28m^2

OpenStudy (anonymous):

20*49 = 980 factors of 980 = 2, 4, 5, 7, 10, 14, 20, 28, 35, 49, 70, 98, 140, 196, 245, 490 choose any 2 at reverse end pairs (2,490),(4,196),(ect) total of 4,196 = 200. Split 200 into 2 parts based on total. coeficient a = (28/200)*4 coeficient b = (28/200)*196 20x^2 + 0.56x + 27.44x + 49 I'm not sure if you can directly factor this...